In a closed economy, private saving is smaller than investment if government spending exceeds tax revenue.

Select one:

True

False

If there is a surplus of loanable funds, then neither curve shifts, but the quantity of loanable funds supplied increases and the quantity demanded decreases as the interest rate rises to equilibrium.

Select one:

True

False

An increase in the budget deficit would cause a shortage of loanable funds at the original interest rate, which would lead to falling interest rates and a decrease in investment

Select one:

True

False

Credit risk refers to a bond’s probability of default and because they are low risk and attractive to buyers, U.S. government bonds usually pay a low rate of interest.

Select one:

True

False

Suppose that in a closed economy the value of GDP is equal to $64 billion. Taxes are equal to $16 billion, consumption equals $40 billion, government expenditures equal $12 billion and public saving equals $4 billion. The amount of national saving is $12 billion.

Select one:

True

False
Answer #1

(1) False

Y = C + I + G - T + X - M, with usual notation

Y - C - I + T - G = 0 [since X = M = 0 in closed economy]

S - I + T - G = 0 [since S = Y - C]

S - I = G - T

So, if G > T, it means S > I.

(2) False

A surplus arises when quantity supplied (Qs) > quantity demanded (Qd) at a higher-than-equilibrium interest rate. As a result, Qs starts decreasing and Qd starts increasing when interest rate falls and reaches equilibrium interest rate.

(3) False

An increase in the budget deficit will increase demand for loanable funds, which, by increasing interest rate, will cause a shortage of loanable funds at the original interest rate, which would lead to rising interest rates and a decrease in investment.

(4) True

Credit (default) risk indicates the risk of default, making risk-free government bonds sell at higher price, lowering their interest rate.

(5) False

Nation saving ($B) = Private saving + Public saving = (GDP - C) + (T - G) = 64 - 40 + 16 - 12 = 28